# ELLA: Efficient Lifelong Learning for Adapters in Large Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable generalization capabilities across a wide range of downstream tasks, largely attributed to their large-scale pretraining on diverse corpora brown2020language; touvron2023llama; achiam2023gpt. As LLMs are increasingly deployed in real-world applications, they must be continuously adapted to evolving user needs and task distributions for long-term practical deployment. This setting, commonly studied under the paradigm of continual learning (CL), requires models to acquire new knowledge sequentially without costly full-retraining ruvolo2013ella. However, sequential finetuning of LLMs remains highly susceptible to catastrophic forgetting (CF) - the tendency to overwrite prior knowledge when new tasks are introduced mccloskey1989catastrophic, and loss of plasticity - deterioration in the ability to learn new information over time dohare2021continual. These challenges are especially pronounced in rehearsal-free settings, where previously seen data cannot be stored due to privacy or storage constraints chaudhry2019continual.
A promising line of work leverages parameter-efficient fine-tuning (PEFT) strategies, such as Low-Rank Adaptation (LoRA) hu2022lora, to reduce the overhead of task-specific adaptation. Adapter-based approaches restrict updates to a small set of trainable parameters while leaving the base model frozen, making them a natural fit for lower-compute CL qin2021lfpt5; song2023conpet. Yet, even with such modularity, sequential adapter training often yields severe forgetting when past knowledge is not revisited wang2024inscl. Prior efforts to address this include expanding model capacity wang2024rehearsal, isolating weights to reduce interference aljundi2017expert; li2019learn; wang2023rehearsal, subspace orthogonality wang2023orthogonal; liao2025data, or gradient projection from previous tasks qiao2024learn. While effective to varying degrees, these often either limit forward knowledge transfer from past tasks, add storage overhead, or overlook activation-level interference, where forgetting actually manifests ke2021achieving.
A truly effective continual learning framework should balance knowledge retention with representation reuse, preventing harmful interference while allowing useful subspaces to be shared for new tasks wu2024continual. In practice, not all overlap is harmful – low-magnitude directions from past tasks may encode generic linguistic or semantic patterns that can accelerate learning on future tasks if safely reused. Unfortunately, existing CL methods either eliminate all overlap or fuse past knowledge through heavyweight mechanisms like controller networks or rank-conditioned fusion, which limits scalability and increases complexity zhao2024sapt; liu2023vida; liao2025data.
@@
Through extensive experiments on the Standard CL Benchmark zhang2015character, Long Sequence Benchmark razdaibiedina2023progressive and TRACE wang2023trace, we demonstrate that ELLA achieves state-of-the-art performance while scaling effectively across model sizes (from $770M$ to $8B$). Furthermore, ELLA generalizes well across architecture families (e.g., T5 raffel2020exploring, LLaMA touvron2023llama), and unlike prior work, uniquely improves generalization to unseen tasks. ELLA does not rely on task identities during inference, and is hence naturally compatible with the instruction-tuning paradigm wang2022task, thereby preserving the generalization capabilities of LLMs in zero-shot and open-ended settings. Notably, our method can also be seamlessly integrated with existing continual learning methods to further enhance their effectiveness, without requiring additional supervision or auxiliary components. Moreover, ELLA is architecture-agnostic, requires no access to past data, and introduces negligible computational or memory overhead. In summary, our contributions are as follows:
@@ 2 Related Works
Continual Learning   The primary challenge in CL is to balance model stability (resisting catastrophic forgetting) with plasticity (acquiring new knowledge). Prior approaches to address this problem can be broadly categorized by their core mechanism, each presenting a distinct set of trade-offs. (i) Rehearsal-based methods store a subset of past data and interleave it during subsequent training phases. These include strategies such as experience replay riemer2018learning or constrained optimization aljundi2017expert; chaudhry2019continual; he2024seekr that jointly train on current and previous samples. While empirically strong, this approach is often impractical due to significant storage overhead and potential violations of data privacy constraints. (ii) Regularization-based methods add penalty terms to the loss that restrict updates to parameters critical for old tasks du2024unlocking; li2017learning; de2019episodic. For instance, kirkpatrick2017overcoming slows updates on important weights, while farajtabar2020orthogonal constrains new updates to be orthogonal to gradients of old tasks. These methods depend on importance metrics that are often brittle and prevent any forward transfer, limiting adaptability to new tasks. (iii) Architecture-based methods reduce interference through task-specific modules or by expanding model size li2019learn; wang2023rehearsal. While razdaibiedina2023progressive appends a new learned soft prompt per task,  qin2021lfpt5 utilizes a large soft prompt that is continuously trained on all tasks. While this effectively isolates task knowledge, it typically results in a linear growth in parameter count with the number of tasks, posing significant scalability challenges and often requiring explicit task labels at inference.
@@
Parameter-Efficient Continual Learning   Parameter-Efficient Fine-Tuning (PEFT) has emerged as a highly promising substrate for CL, with Low-Rank Adaptation (LoRA) hu2022lora being a particularly effective technique. However, naïvely applying LoRA sequentially leads to significant forgetting owing to interference across tasks. As tasks arrive sequentially, the adapter weights $A_{t}$, $B_{t}$ learned for a new task $t$ are trained from scratch without awareness of previous tasks’ LoRA subspaces. Consequently, overlapping update directions in parameter space may inadvertently override past knowledge, especially under limited parameter budgets. Without coordination across tasks, these LoRA components can destructively interfere in spaces crucial for previous tasks, degrading performance on earlier tasks while optimizing for the current one. Recent works in CL literature addressing this issue primarily falls into two schools of thought.
@@
The first approach applies strict orthogonality to enforce zero overlap between the LoRA subspaces of different tasks wang2023orthogonal; qiao2024learn. While this hard constraint effectively mitigates interference, it imposes an overly restrictive inductive bias. By precluding any form of representation sharing, it severely inhibits beneficial forward transfer and prevents reuse of low-importance shared space across related tasks. Worse still, as tasks accumulate, the available orthogonal space quickly diminishes, leading to progressively less capacity for future learning. This rigid decoupling causes inefficient use of adapter capacity, hindering the goal of cumulative learning. Furthermore, storing all past LoRA weights incurs high memory overhead that scales linearly with the number of tasks, further limiting scalability.
@@
The second approach involves more complex fusion and replay mechanisms. Methods like DATA liao2025data or Recurrent KIF feng2025recurrent often rely on replay buffers, multiple LoRA adapters per task, or intricate gating mechanisms to learn how to compose adapters. These methods reintroduce significant architectural complexity and computational overhead, sacrificing the very simplicity and efficiency that make PEFT an attractive paradigm for CL in the first place.
@@ 3.1 Problem Formulation
Low-Rank Adaptation (LoRA). Our method builds on LoRA hu2022lora, a PEFT technique that leverages the observation that fine-tuning large pre-trained models (PTMs) can be effectively performed within a low-dimensional subspace aghajanyan2020intrinsic. Given a frozen pre-trained weight matrix $W_{\mathrm{init}}\in\mathbb{R}^{d\times k}$, LoRA introduces a low-rank update $\Delta W=AB$, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ with $r\ll\min(d,k)$. The original weights $W_{\mathrm{init}}$ remain unchanged during training, and only the low-rank matrices $A$ and $B$ are optimized. During the forward pass, the original operation $h=W_{\mathrm{init}}x$ becomes:
@@
This design reduces memory and compute overhead significantly, while achieving competitive performance to full fine-tuning across model families and benchmarks houlsby2019parameter.
@@ 3.2 Subspace-Aware Continual Adaptation
By construction, $\mathcal{W}_{\text{past}}$ encodes the dominant directions heavily used by prior tasks. Motivated by the observation that high-magnitude LoRA components are typically more task-discriminative aghajanyan2020intrinsic, we introduce the ELLA alignment penalty as follows:
@@ 4.1 Datasets
Standard CL Benchmark (SC) is a popular CL benchmark designed for language models, comprising five text classification datasets introduced by zhang2015character. Following wang2023orthogonal, we select AG News, Amazon Reviews, DBpedia, and Yahoo Answers, and form three shuffled task sequences, denoted as Orders $1$, 22, and $3$.
Long Sequence Benchmark (LS) extends the standard CL setup by incorporating $15$ tasks, including five classification datasets, nine GLUE and SuperGLUE benchmarks, and the IMDB dataset razdaibiedina2023progressive. Following the protocol of prior works liao2025data, we train each task using $1,000$ randomly selected samples, with $500$ samples per class reserved for testing. These tasks are also arranged into shuffled sequences Orders $4$, 55, and $6$.
TRACE is a benchmark tailored for CL in LLMs, covering a diverse set of $8$ tasks that span domains such as multiple-choice question answering, multilingual understanding, code generation, and mathematical reasoning wang2023trace. See App. B.2 for details on tasks and orderings.
@@ 4.2 Metrics
Let $a_{i,j}$ denote the testing performance on the $j$-th task after training on the $i$-th task. We evaluate across: Overall Accuracy (OA) chaudhry2018riemannian: The average accuracy across all tasks after training on the last task, i.e., $\text{OA}_{\mathcal{T}}=\frac{1}{\mathcal{T}}\sum_{t=1}^{\mathcal{T}}a_{\mathcal{T},t}$; Forward Transfer (FWT) lopez2017gradient: measures how much knowledge from previous tasks transfers to a new task, i.e., $\text{FWT}_{\mathcal{T}}=\frac{1}{\mathcal{T}}\sum_{t=1}^{\mathcal{T}}(a_{t,t}-a_{0,t})$, where $a_{0,t}$ refers to the performance of training task $t$ individually; Backward Transfer (BWT) ke2022continual: measures how much the learning of subsequent tasks influences the performance of prior tasks, i.e., $\text{BWT}_{\mathcal{T}}=\frac{1}{\mathcal{T}-1}\sum_{t=1}^{\mathcal{T}-1}(a_{\mathcal{T},t}-a_{t,t})$. Moreover, we also report general ability (GA) – the average generalization performance across unseen datasets post CL, and delta general ability (DeltaGA) – change in GA relative to the original LLM.
@@ 4.3 Baselines
We compare ELLA against a comprehensive suite of CL baselines. This includes naïve sequential fine-tuning approaches (SeqFT, SeqLoRA), classic regularization methods (EWC kirkpatrick2017overcoming, LwF li2017learning), and modern prompt-tuning techniques (L2P wang2022learning, LFPT5 qin2021lfpt5). We also benchmark against state-of-the-art adapter-based methods, including those that enforce strict orthogonality (O-LoRA wang2023orthogonal, LB-CL qiao2024learn) as well as those that combine orthogonality with complex decomposition or replay mechanisms (DATA liao2025data, Recurrent KIF feng2025recurrent). Details of baselines are provided in Appendix B.3.
@@
T5-Large SeqFT de2019episodic 18.9 24.9 41.7 28.5 7.4 7.3 7.4 7.4 -
@@
EWC kirkpatrick2017overcoming 46.3 45.3 52.1 47.9 44.9 44.0 45.4 44.8 -
LwF li2017learning 52.7 52.9 48.4 51.3 49.7 42.8 46.9 46.5 -
L2P wang2022learning 59.0 60.5 59.9 59.8 57.7 53.6 56.6 56.0 -
@@
LB-CL qiao2024learn 76.9 76.5 76.8 76.7 68.4 67.3 71.8 69.2 -
O-LoRA wang2023orthogonal 73.5 71.4 70.0 71.6 65.4 65.2 65.2 65.3 23.1
   + MIGU du2024unlocking 77.1 77.0 75.6 76.6 67.3 68.5 74.0 70.0 -
DATA liao2025data 71.5 70.5 68.0 70.0 71.5 70.5 68.0 70.0 16.7
@@
O-LoRA wang2023orthogonal 69.39 67.58 71.44 69.46 69.54 64.42 66.50 66.82 28.45
DATA liao2025data 76.10 75.69 75.52 75.77 71.34 70.77 72.95 71.69 31.33
@@ 4.4 Implementation Details
We evaluate ELLA on a range of LLMs including encoder-decoder T55 models raffel2020exploring (T55-Large, T55-XL), and decoder-only LLaMA models touvron2023llama (LLaMA-$3.1$ 88B). ELLA training is performed with the DeepSpeed library using V100100 16GB GPUs for T5 and A4040 48GB GPUs for LLaMA, while larger baselines such as DATA liao2025data require H100100 80GB GPUs. All experiments are performed with instruction tuning wei2021finetuned. Following wang2023orthogonal, we apply LoRA modules (rank=$8$) to the Attention Q-V layers and report the average result of $3$ runs. More details in App. B.6.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable generalization capabilities across a wide range of downstream tasks, largely attributed to their large-scale pretraining on diverse corpora Brown et al. (2020); Touvron et al. (2023); Achiam et al. (2023). As LLMs are increasingly deployed in real-world applications, they must be continuously adapted to evolving user needs and task distributions for long-term practical deployment. This setting, commonly studied under the paradigm of continual learning (CL), requires models to acquire new knowledge sequentially without costly full-retraining Ruvolo and Eaton (2013). However, sequential finetuning of LLMs remains highly susceptible to catastrophic forgetting (CF) - the tendency to overwrite prior knowledge when new tasks are introduced McCloskey and Cohen (1989), and loss of plasticity - deterioration in the ability to learn new information over time Dohare et al. (2021). These challenges are especially pronounced in rehearsal-free settings, where previously seen data cannot be stored due to privacy or storage constraints Chaudhry et al. (2019).
A promising line of work leverages parameter-efficient fine-tuning (PEFT) strategies, such as Low-Rank Adaptation (LoRA) Hu et al. (2022), to reduce the overhead of task-specific adaptation. Adapter-based approaches restrict updates to a small set of trainable parameters while leaving the base model frozen, making them a natural fit for lower-compute CL Qin and Joty (2021); Song et al. (2023). Yet, even with such modularity, sequential adapter training often yields severe forgetting when past knowledge is not revisited Wang et al. (2024b). Prior efforts to address this include expanding model capacity Wang et al. (2024a), isolating weights to reduce interference Aljundi et al. (2017); Li et al. (2019); Wang et al. (2023c), subspace orthogonality Wang et al. (2023a); Liao et al. (2025), or gradient projection from previous tasks Qiao and Mahdavi (2024). While effective to varying degrees, these often either limit forward knowledge transfer from past tasks, add storage overhead, or overlook activation-level interference, where forgetting actually manifests Ke et al. (2021).
A truly effective continual learning framework should balance knowledge retention with representation reuse, preventing harmful interference while allowing useful subspaces to be shared for new tasks Wu et al. (2024). In practice, not all overlap is harmful – low-magnitude directions from past tasks may encode generic linguistic or semantic patterns that can accelerate learning on future tasks if safely reused. Unfortunately, existing CL methods either eliminate all overlap or fuse past knowledge through heavyweight mechanisms like controller networks or rank-conditioned fusion, which limits scalability and increases complexity Zhao et al. (2024); Liu et al. (2023); Liao et al. (2025).
@@
Through extensive experiments on the Standard CL Benchmark Zhang et al. (2015), Long Sequence Benchmark Razdaibiedina et al. (2023) and TRACE Wang et al. (2023b), we demonstrate that ELLA achieves state-of-the-art performance while scaling effectively across model sizes (from $770M$ to $8B$). Furthermore, ELLA generalizes well across architecture families (e.g., T5 Raffel et al. (2020), LLaMA Touvron et al. (2023)), and unlike prior work, uniquely improves generalization to unseen tasks. ELLA does not rely on task identities during inference, and is hence naturally compatible with the instruction-tuning paradigm Wang et al. (2022a), thereby preserving the generalization capabilities of LLMs in zero-shot and open-ended settings. Notably, our method can also be seamlessly integrated with existing continual learning methods to further enhance their effectiveness, without requiring additional supervision or auxiliary components. Moreover, ELLA is architecture-agnostic, requires no access to past data, and introduces negligible computational or memory overhead. In summary, our contributions are as follows:
@@ 2 Related Works
Continual Learning   The primary challenge in CL is to balance model stability (resisting catastrophic forgetting) with plasticity (acquiring new knowledge). Prior approaches to address this problem can be broadly categorized by their core mechanism, each presenting a distinct set of trade-offs. (i) Rehearsal-based methods store a subset of past data and interleave it during subsequent training phases. These include strategies such as experience replay Riemer et al. (2018) or constrained optimization Aljundi et al. (2017); Chaudhry et al. (2019); He et al. (2024) that jointly train on current and previous samples. While empirically strong, this approach is often impractical due to significant storage overhead and potential violations of data privacy constraints. (ii) Regularization-based methods add penalty terms to the loss that restrict updates to parameters critical for old tasks Du et al. (2024); Li and Hoiem (2017); de Masson D’Autume et al. (2019). For instance, Kirkpatrick et al. (2017) slows updates on important weights, while Farajtabar et al. (2020) constrains new updates to be orthogonal to gradients of old tasks. These methods depend on importance metrics that are often brittle and prevent any forward transfer, limiting adaptability to new tasks. (iii) Architecture-based methods reduce interference through task-specific modules or by expanding model size Li et al. (2019); Wang et al. (2023c). While Razdaibiedina et al. (2023) appends a new learned soft prompt per task,  Qin and Joty (2021) utilizes a large soft prompt that is continuously trained on all tasks. While this effectively isolates task knowledge, it typically results in a linear growth in parameter count with the number of tasks, posing significant scalability challenges and often requiring explicit task labels at inference.
@@
Parameter-Efficient Continual Learning   Parameter-Efficient Fine-Tuning (PEFT) has emerged as a highly promising substrate for CL, with Low-Rank Adaptation (LoRA) Hu et al. (2022) being a particularly effective technique. However, naïvely applying LoRA sequentially leads to significant forgetting owing to interference across tasks. As tasks arrive sequentially, the adapter weights $A_{t}$, $B_{t}$ learned for a new task $t$ are trained from scratch without awareness of previous tasks’ LoRA subspaces. Consequently, overlapping update directions in parameter space may inadvertently override past knowledge, especially under limited parameter budgets. Without coordination across tasks, these LoRA components can destructively interfere in spaces crucial for previous tasks, degrading performance on earlier tasks while optimizing for the current one. Recent works in CL literature addressing this issue primarily falls into two schools of thought.
@@
The first approach applies strict orthogonality to enforce zero overlap between the LoRA subspaces of different tasks Wang et al. (2023a); Qiao and Mahdavi (2024). While this hard constraint effectively mitigates interference, it imposes an overly restrictive inductive bias. By precluding any form of representation sharing, it severely inhibits beneficial forward transfer and prevents reuse of low-importance shared space across related tasks. Worse still, as tasks accumulate, the available orthogonal space quickly diminishes, leading to progressively less capacity for future learning. This rigid decoupling causes inefficient use of adapter capacity, hindering the goal of cumulative learning. Furthermore, storing all past LoRA weights incurs high memory overhead that scales linearly with the number of tasks, further limiting scalability.
@@
The second approach involves more complex fusion and replay mechanisms. Methods like DATA Liao et al. (2025) or Recurrent KIF Feng et al. (2025) often rely on replay buffers, multiple LoRA adapters per task, or intricate gating mechanisms to learn how to compose adapters. These methods reintroduce significant architectural complexity and computational overhead, sacrificing the very simplicity and efficiency that make PEFT an attractive paradigm for CL in the first place.
@@ 3.1 Problem Formulation
Low-Rank Adaptation (LoRA). Our method builds on LoRA Hu et al. (2022), a PEFT technique that leverages the observation that fine-tuning large pre-trained models (PTMs) can be effectively performed within a low-dimensional subspace Aghajanyan et al. (2020). Given a frozen pre-trained weight matrix $W_{\mathrm{init}}\in\mathbb{R}^{d\times k}$, LoRA introduces a low-rank update $\Delta W=AB$, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ with $r\ll\min(d,k)$. The original weights $W_{\mathrm{init}}$ remain unchanged during training, and only the low-rank matrices $A$ and $B$ are optimized. During the forward pass, the original operation $h=W_{\mathrm{init}}x$ becomes:
@@
This design reduces memory and compute overhead significantly, while achieving competitive performance to full fine-tuning across model families and benchmarks Houlsby et al. (2019).
@@ 3.2 Subspace-Aware Continual Adaptation
By construction, $\mathcal{W}_{\text{past}}$ encodes the dominant directions heavily used by prior tasks. Motivated by the observation that high-magnitude LoRA components are typically more task-discriminative Aghajanyan et al. (2020), we introduce the ELLA alignment penalty as follows:
@@ 4.1 Datasets
Standard CL Benchmark (SC) is a popular CL benchmark designed for language models, comprising five text classification datasets introduced by Zhang et al. (2015). Following Wang et al. (2023a), we select AG News, Amazon Reviews, DBpedia, and Yahoo Answers, and form three shuffled task sequences, denoted as Orders $1$, 22, and $3$.
Long Sequence Benchmark (LS) extends the standard CL setup by incorporating $15$ tasks, including five classification datasets, nine GLUE and SuperGLUE benchmarks, and the IMDB dataset Razdaibiedina et al. (2023). Following the protocol of prior works Liao et al. (2025), we train each task using $1,000$ randomly selected samples, with $500$ samples per class reserved for testing. These tasks are also arranged into shuffled sequences Orders $4$, 55, and $6$.
TRACE is a benchmark tailored for CL in LLMs, covering a diverse set of $8$ tasks that span domains such as multiple-choice question answering, multilingual understanding, code generation, and mathematical reasoning Wang et al. (2023b). See App. B.2 for details on tasks and orderings.
@@ 4.2 Metrics
Let $a_{i,j}$ denote the testing performance on the $j$-th task after training on the $i$-th task. We evaluate across: Overall Accuracy (OA) Chaudhry et al. (2018): The average accuracy across all tasks after training on the last task, i.e., $\text{OA}_{\mathcal{T}}=\frac{1}{\mathcal{T}}\sum_{t=1}^{\mathcal{T}}a_{\mathcal{T},t}$; Forward Transfer (FWT) Lopez-Paz and Ranzato (2017): measures how much knowledge from previous tasks transfers to a new task, i.e., $\text{FWT}_{\mathcal{T}}=\frac{1}{\mathcal{T}}\sum_{t=1}^{\mathcal{T}}(a_{t,t}-a_{0,t})$, where $a_{0,t}$ refers to the performance of training task $t$ individually; Backward Transfer (BWT) Ke and Liu (2022): measures how much the learning of subsequent tasks influences the performance of prior tasks, i.e., $\text{BWT}_{\mathcal{T}}=\frac{1}{\mathcal{T}-1}\sum_{t=1}^{\mathcal{T}-1}(a_{\mathcal{T},t}-a_{t,t})$. Moreover, we also report general ability (GA) – the average generalization performance across unseen datasets post CL, and delta general ability (DeltaGA) – change in GA relative to the original LLM.
@@ 4.3 Baselines
We compare ELLA against a comprehensive suite of CL baselines. This includes naïve sequential fine-tuning approaches (SeqFT, SeqLoRA), classic regularization methods (EWC Kirkpatrick et al. (2017), LwF Li and Hoiem (2017)), and modern prompt-tuning techniques (L2P Wang et al. (2022b), LFPT5 Qin and Joty (2021)). We also benchmark against state-of-the-art adapter-based methods, including those that enforce strict orthogonality (O-LoRA Wang et al. (2023a), LB-CL Qiao and Mahdavi (2024)) as well as those that combine orthogonality with complex decomposition or replay mechanisms (DATA Liao et al. (2025), Recurrent KIF Feng et al. (2025)). Details of baselines are provided in Appendix B.3.
@@
T5-Large SeqFT de Masson D’Autume et al. (2019) 18.9 24.9 41.7 28.5 7.4 7.3 7.4 7.4 -
@@
EWC Kirkpatrick et al. (2017) 46.3 45.3 52.1 47.9 44.9 44.0 45.4 44.8 -
LwF Li and Hoiem (2017) 52.7 52.9 48.4 51.3 49.7 42.8 46.9 46.5 -
L2P Wang et al. (2022b) 59.0 60.5 59.9 59.8 57.7 53.6 56.6 56.0 -
@@
LB-CL Qiao and Mahdavi (2024) 76.9 76.5 76.8 76.7 68.4 67.3 71.8 69.2 -
O-LoRA Wang et al. (2023a) 73.5 71.4 70.0 71.6 65.4 65.2 65.2 65.3 23.1
   + MIGU Du et al. (2024) 77.1 77.0 75.6 76.6 67.3 68.5 74.0 70.0 -
DATA Liao et al. (2025) 71.5 70.5 68.0 70.0 71.5 70.5 68.0 70.0 16.7
@@
O-LoRA Wang et al. (2023a) 69.39 67.58 71.44 69.46 69.54 64.42 66.50 66.82 28.45
DATA Liao et al. (2025) 76.10 75.69 75.52 75.77 71.34 70.77 72.95 71.69 31.33
@@ 4.4 Implementation Details
We evaluate ELLA on a range of LLMs including encoder-decoder T55 models Raffel et al. (2020) (T55-Large, T55-XL), and decoder-only LLaMA models Touvron et al. (2023) (LLaMA-$3.1$ 88B). ELLA training is performed with the DeepSpeed library using V100100 16GB GPUs for T5 and A4040 48GB GPUs for LLaMA, while larger baselines such as DATA Liao et al. (2025) require H100100 80GB GPUs. All experiments are performed with instruction tuning Wei et al. (2021). Following Wang et al. (2023a), we apply LoRA modules (rank=$8$) to the Attention Q-V layers and report the average result of $3$ runs. More details in App. B.6.